OpenStudy (faman39):

During the winter, an ice cream shop has 60% of the employees it has during the summer. Which equation shows that during the winter, the ice cream shop has 9 employees? 60 = 0.09 *s w = 0.6 *9 9 = 0.6*s w = 9% * 60

OpenStudy (anonymous):

9 = 0.6*s

OpenStudy (anonymous):

60% 0f summer employee=9 (60/100)s=9 0.6 s=9

OpenStudy (anonymous):

any question?

OpenStudy (faman39):

No, so it will be 9 = 0.6 s

OpenStudy (anonymous):

yup

OpenStudy (faman39):

Thank you soo much! :)

OpenStudy (anonymous):

YW, good luck now, have fun now...lol :D

OpenStudy (faman39):

:) thank you